A bucket that weighs 5 lb and a rope of negligible weight are used to draw water from a well that is 80 ft deep. The bucket is filled with 36 lb of water and is pulled up at a rate of 2 ft/s, but water leaks out of a hole in the bucket at a rate of 0.2 lb/s. Find the work done in pulling the bucket to the top of the well. Show how to approximate the required work by a Riemann sum. (Let x be the height in feet above the bottom of the well. Enter xi* as xi.)
Answer #1

Given: wt. of the buclect = 5 lb Depth of well .=80 ft water in the buclect = 36 lb Rate of pulling aft/ see . Rate of healea(2) For water: Rate of pulling = 2ft/ see i in second, water moves for 2 feet is water moves to 80 ft = 80 = 40 see water movkLinear equation of a line is y = matc -) weight = ma + weight = 16x +C When weight = 28 lb, we have a=0. ... 28 = 1 (0)+Ć 3)Hence, Total Work done = 400+2360 = 2760

Hence work done = 2760 ft-lb